Il ela + Ek | / 2 ) تا (3- = =ة
Answer #1

as (1,-3)<b(-1.2) Reles &B-2. <a (3) = (31) bb. (-2-1-22) (21) x + e (3+2 M-4) x (Strs) -- || aa + BE || = JST+(-) 2x+25 So e